Calculate the work done by a 14N force at 40° to the horizontal to move a 505g box a horizontal distance of 40m across a rough floor at a constant speed of 0.5m/s.
1
Expert's answer
2020-03-23T13:05:36-0400

"F=14" "N"

"\alpha=40^o"

"m=505" "g"

"S=40" "m"

"v=0.5" "m\/s"


Answer:

"A=FScos(\alpha)=14*40*cos(40^o)=498.985" "J"